# Networking Systems for Video Anomaly Detection: A Tutorial and Survey

### v2

@@ Abstract.
The increasing utilization of surveillance cameras in smart cities, coupled with the surge of online video applications, has heightened concerns regarding public security and privacy protection, which propelled automated Video Anomaly Detection (VAD) into a fundamental research task within the Artificial Intelligence (AI) community. With the advancements in deep learning and edge computing, VAD has made significant progress and advances synergized with emerging applications in smart cities and video internet, which has moved beyond the conventional research scope of algorithm engineering to deployable Networking Systems for VAD (NSVAD), a practical hotspot for intersection exploration in the AI, IoVT, and computing fields. In this article, we delineate the foundational assumptions, learning frameworks, and applicable scenarios of various deep learning-driven VAD routes, offering an exhaustive tutorial for novices in NSVAD. This article elucidates core concepts by reviewing recent advances and typical solutions and aggregating available research resources accessible at https://github.com/fdjingliu/NSVAD. Additionally, we showcase our latest NSVAD research in industrial IoT and smart cities, along with an end-cloud collaborative architecture for deployable NSVAD. Lastly, this article projects future development trends and discusses how the integration of AI and computing technologies can address existing research challenges and promote open opportunities, serving as an insightful guide for prospective researchers and engineers.
@@ 1. Introduction
As one of the core technologies of the ubiquitous Internet of Video Things (IoVT), Video Anomaly Detection (VAD) aims to use video sensors to automatically discover unexpected spatial-temporal patterns and detect unusual events that may cause security problems or economic losses, such as traffic accidents, violent behaviors, and offending contents (Liu et al., 2024d). With the widespread use of surveillance cameras in smart cities (Ramachandra et al., 2020) and the boom of online video applications powered by 4/5G communication technologies, traditional human inspection is no longer able to accurately monitor the video data generated around the clock, which is not only time-consuming and labor-intensive but also poses the risk of leaking important information (e.g., biometrics and sensitive speech). In contrast, VAD-empowered IoVT applications (Jedari et al., 2020), such as Intelligent Surveillance Systems (IVSS) and automated content analysis platforms, can process massive video streams online and detect events of interest in real-time, sending only noteworthy anomaly parts for human review, significantly reducing data storage and communication costs, and helping to eliminate public concerns about data security and privacy protection. As a result, VAD has gained widespread attention in academia and industry over the last decade and has been used in emerging fields such as information forensics (Yu et al., 2023), industrial manufacturing (Liu et al., 2024a) in smart cities as well as online content analysis in mobile video applications (Yin et al., 2024).
@@
VAD extends the data scope of conventional Anomaly Detection (AD) from time series, images, and graphs to video, which not only needs to cope with the endogenous data complexity, but also needs to take into account the computational and communication costs in resource-limited devices (Jiang et al., 2021). Specifically, the inherent high-dimensional structure of video data, high information density and redundancy, heterogeneity of temporal and spatial patterns, and feature entanglement between foreground targets and background scenes make VAD more challenging than traditional AD tasks at the levels of representation learning and anomaly discrimination (Mishra et al., 2024). Existing studies (Liu et al., 2018b, 2024b; Bao et al., 2022; Li et al., 2024a) have shown that high-performance VAD models need to target the modeling of appearance and motion information, i.e., the difference between regular events and anomalous examples in both spatial and temporal dimensions. In contrast to time series AD which mainly measures periodic temporal patterns of variables, and image AD which only focuses on spatial contextual deviations, VAD needs to extract both discriminative spatial and temporal features from a large amount of redundant information (e.g., repetitive temporal contexts and label-independent data distributions), as well as to learn the differences between normal and anomalous events in terms of their local appearances and global motions (Qiu et al., 2024).
@@
However, video anomalies are ambiguous and subjective (Hojjati et al., 2024). The same behavior may be categorized as different types in different scenarios and settings. For example, riding a horse in a grassland is usually normal, whereas a horse appearing on a highway would be considered an anomaly. On the one hand, compared to regular events, anomalies in the real world are difficult to be comprehensively predefined and have a much lower overall frequency of occurrence, making them difficult to collect. Labeling a sufficient number of all possible abnormal samples for model training is almost impossible. As a result, traditional supervised learning-based classification models are usually ineffective in dealing with AD tasks (Mašková et al., 2024). On the other hand, since video storage and transmission costs are significantly higher than other data modalities, engineers favor processing such data on the end or edge side to reduce communication overhead. As we all know, such devices, including surveillance cameras, smartphones, and local servers, are computationally resource-limited. Therefore, it is highly practical to develop deployable VAD systems for real-world applications, which requires concerted efforts by AI beyond communities.
@@
In this article, we extend the conventional scope of VAD from algorithm engineering on spatial-temporal anomaly detection to practical research towards real-world applications, termed Networking Systems for Video Anomaly Detection (NSVAD), to engage a broader readership from the IoT and computing communities. According to research objectives and involved domains, NSVAD is delineated into the hierarchical architecture shown in Fig. 1, encompassing: 1) Hardware Layer, consisting of various video sensors, communication units, and computing servers, etc, responsible for data acquisition, transmission, processing, and result reporting, as well as device networking; 2) System Layer, targeting resource optimization and algorithm deployment platforms for large-scale IoT applications, linking the Hardware Layer and the Algorithm Layer, supporting the configurable deployment of VAD tasks on various terminals; 3) Algorithm Layer, focusing on the development of detection algorithms and scene-specific models driven by artificial intelligence, especially deep learning; 4) Application Layer, encompassing IVSS in modern factories, agriculture, and smart cities, as well as various online video applications powered mobile internet. Most existing works belong to the Algorihm layer and solely concentrate on VAD model design, overlooking resource costs and challenges in real-world scenarios. For large-scale IoT and mobile video internet, NSVAD with stable performance and reasonable overheads that support online detection necessitates sensor networking research and support for resource optimization from computing communities. We review the recent advancements and typical methods in the algorithm layer and provide our latest explorations in the system layer to inspire readers to develop NSVAD toward real-world scenarios.
@@
Thanks to the development of edge AI (Walia et al., 2023) and artificial neural networks (Li et al., 2021b), deep learning-driven NSVAD algorithms have made significant progress in recent years and derived Unsupervised (UVAD), Weakly-supervised (WsVAD), and Fully-unsupervised (FuVAD) routes (Liu et al., 2024d). They have liberated human beings from massive videos analysis works and alleviated public information security concerns. Compared to early manual feature engineering, deep architectures such as convolutional neural networks and attention mechanisms can extract spatial-temporal representations from video sequences end-to-end without applying human a priori, empowering IVSS to process videos in different resolutions and scenarios. Therefore, researchers in this field are currently focusing on deep structure design and optimization learning strategies. They have creatively proposed multimodal VAD (Wu et al., 2020, 2022; Tian et al., 2021), Open-Set AVD (OSVAD) (Acsintoae et al., 2022; Zhu et al., 2022), Open-Vocabulary VAD (OVVAD) (Wu et al., 2024c), video anomaly segmentation (Tian et al., 2024), and anomaly retrial (Wu et al., 2024b) tasks as well as integrated detection systems that can be deployed in practical scenes, such as modern manufacturing (Liu et al., 2024b), smart city (Liu et al., 2023e), and automated driving (Wang et al., 2021a). In addition to algorithm design, researchers from Networking Systems of AI (i.e., research on the deep convergence of communication and AI) (Song et al., 2022) and IoVT (i.e., subfields of IoT focusing on video sensor design, networking, and data processing) (Chen, 2020) have begun to explore the design deployment-oriented VAD systems to collaboratively deal with multiple challenges (e.g., multi-view cross-scene heterogeneous videos and communication overheads) that come from the dynamic scenarios at the application layer and the limited resources at the hardware layer. These explorations and progress greatly expand the research boundaries and application scenarios of VAD, promoting it as an intelligent system science, i.e., NSVAD.
@@
Although there have been some reviews (Pang et al., 2021; Cook et al., 2019; Santhosh et al., 2020; Li et al., 2023; Samaila et al., 2024) focusing on AD and combing its related work, due to the limited research horizon, the early works usually regard VAD as a fringe research task in the AD community. They focus on time series or images but lack an illuminating analysis of the AD task in video data. Recent survey papers (Nayak et al., 2021; Ramachandra et al., 2020; Chandrakala et al., 2022) continue to focus on unsupervised NSVAD routes in the same vein as the conventional AD task, i.e., using only normal samples to train generative models to learn the prototypical patterns of regular events and to discriminate uncharacterizable test samples as anomalies (Luo et al., 2021). Such reviews, while providing a comprehensive taxonomy of VAD research from the outlier detection perspective, have been informative in the last decade when unsupervised methods have dominated the NSVAD algorithm research. However, they ignore the emerging weakly-supervised (Feng et al., 2021a; Tian et al., 2021) and fully unsupervised (Pang et al., 2020; Zaheer et al., 2022) routes, which are of limited value in guiding further research.
@@
Considering the differences in knowledge bases and orientations of readers, this article provides an in-depth analysis of the basic concepts, related knowledge, and recent advances involved in NSVAD research and summarizes the available research resources. We systematically analyze the assumptions, frameworks, scenarios, advantages, and disadvantages of unsupervised, weakly supervised, and fully unsupervised VAD routes and explain in detail the relevant domain knowledge involved in each route. In addition, we introduce our NSVAD systems designed for dynamic environments in complex scenarios such as industrial IoT and smart cities to guide the NSVAD research in specific applications. Finally, we forecast the research challenges, trends, and possible opportunities to inspire future exploration.
@@ 1.1. Attention Analysis
We searched the number of publications and citations with the topic of ”video anomaly detection” in various mainstream academic databases (e.g., IEEE Xplore, ACM Digital Library, SpringerLink, ScienceDirect, and DBLP) to quantitatively present the research hotness, as shown in Fig. 2(a). Early VAD works were limited by hand-crafted features, which cannot handle complicated videos and rely on human a priori, thus having a slow start. Encouraged by the remarkable success of deep learning in video understanding tasks (e.g., action recognition, scene understanding, and multimodal perception), VAD research saw a boom after 2010, with an explosive growth in the number of publications and citations that continues to the present. On the one hand, the spread of surveillance cameras and streaming media platforms has provided sufficient data support for NSVAD research, making it possible to train large-scale deep models with high-performance GPUs. On the other hand, the increasing demand for offending video content detection in various scenarios drives many researchers and engineers from AI and IoT fields to devote themselves to NSVAD research.
@@ 1.2. Related Work
2021 (Nayak et al., 2021) Deep learning-besdUnsupervised VAD
@@
2021 (Pang et al., 2021) Deep learing-besd Unsupervised AD
@@
2024 (Liu et al., 2024d) Generalized VAD
2024 Ours NSVAD routes and cases in video IoT
@@
To our knowledge, this is the first tutorial-type paper on NSVAD, providing a systematic overview of the basics, recent advances, and practical applications of various VAD routes. Previous papers (Liu et al., 2024d; Ramachandra et al., 2020; Chandrakala et al., 2022; Kiran et al., 2018) primarily focus on literature reviews, while conference tutorials lack systematic content, as shown in Table 1. Considering non-specialists’ limited background knowledge, this article emphasizes clear explanations of basic concepts and models. We introduce task definitions, learning frameworks, and practical cases for unsupervised, weakly supervised, and fully unsupervised VAD, as well as emerging tasks like open-set (Zhu et al., 2022), open vocabulary (Wu et al., 2024c), and glance VAD (Zhang et al., 2024b).
Initially, VAD was considered a fringe topic in the broader AD community and only briefly mentioned in AD overviews (Cook et al., 2019; Pang et al., 2021), lacking comprehensive surveys. Recent efforts have started organizing the state of VAD research (Ramachandra et al., 2020; Chandrakala et al., 2022; Raja et al., 2022; Nayak et al., 2024), but they focus mainly on unsupervised methods and overlook rising research routes like WsVAD (Sultani et al., 2018) and FuVAD (Wu et al., 2020). These routes offer reliable performance in real-world applications, and their importance is increasingly recognized. Fully unsupervised methods, in particular, allow efficient VAD model learning from large-scale video streams. Recent works, such as (Liu et al., 2024d), summarize these developments but do not cover emerging tasks like OSVAD (Acsintoae et al., 2022; Zhu et al., 2022) and OVVAD (Wu et al., 2024c), which are of crucial value for IoT applications.
Researchers in the AD community have paid attention to the continued progress and application prospects of VAD and introduced it to attendees through conference tutorials. For example, Pang et al. organized a tutorial titled ”Recent Advances in Anomaly Detection” at CVPR2023, focusing on recent work on deep learning-driven unsupervised and weakly supervised VAD. We think this is a good start and highly appreciate the contribution of the organizers to the field. However, a more unified conceptual statement, more comprehensive documentation, and more systematic analysis of challenges and opportunities are necessary to inspire a wider community of readers.

### v3

@@ Abstract.
The increasing utilization of surveillance cameras in smart cities, coupled with the surge of online video applications, has heightened concerns regarding public security and privacy protection, which propelled automated Video Anomaly Detection (VAD) into a fundamental research task within the Artificial Intelligence (AI) community. With the advancements in deep learning and edge computing, VAD has made significant progress and advances synergized with emerging applications in smart cities and video internet, which has moved beyond the conventional research scope of algorithm engineering to deployable Networking Systems for VAD (NSVAD), a practical hotspot for intersection exploration in the AI, IoVT, and computing fields. In this article, we delineate the foundational assumptions, learning frameworks, and applicable scenarios of various deep learning-driven VAD routes, offering an exhaustive tutorial for novices in NSVAD. In addition, this article elucidates core concepts by reviewing recent advances and typical solutions and aggregating available research resources accessible at https://github.com/fdjingliu/NSVAD. Lastly, this article projects future development trends and discusses how the integration of AI and computing technologies can address existing research challenges and promote open opportunities, serving as an insightful guide for prospective researchers and engineers.
@@ 1. Introduction
As one of the core technologies of the ubiquitous Internet of Video Things (IoVT), Video Anomaly Detection (VAD) aims to use video sensors to automatically discover unexpected spatial-temporal patterns and detect unusual events that may cause security problems or economic losses, such as traffic accidents, violent behaviors, and offending contents (Liu et al., 2024h). With the widespread use of surveillance cameras in smart cities (Ramachandra et al., 2020) and the boom of online video applications powered by 4/5G communication technologies, traditional human inspection is no longer able to accurately monitor the video data generated around the clock, which is not only time-consuming and labor-intensive but also poses the risk of leaking important information (e.g., biometrics and sensitive speech). In contrast, VAD-empowered IoVT applications (Jedari et al., 2020), such as Intelligent Video Surveillance Systems (IVSS) and automated content analysis platforms, can process massive video streams online and detect events of interest in real-time, sending only noteworthy anomaly parts for human review, significantly reducing data storage and communication costs, and helping to eliminate public concerns about data security and privacy protection. As a result, VAD has gained widespread attention in academia and industry over the last decade and has been used in emerging fields (Cao et al., 2024b; Peng et al., 2025; Cao et al., 2025a) such as information forensics (Yu et al., 2023), industrial manufacturing (Liu et al., 2024a; Wang et al., 2024b) in smart cities as well as online content analysis in mobile video applications (Yin et al., 2024).
@@
VAD extends the data scope of conventional Anomaly Detection (AD) from time series, images, and graphs to video, which not only needs to cope with the endogenous data complexity, but also needs to take into account the computational and communication costs in resource-limited devices (Jiang et al., 2021). Specifically, the inherent high-dimensional structure of video data, high information density and redundancy, heterogeneity of temporal and spatial patterns, and feature entanglement between foreground targets and background scenes make VAD more challenging than traditional AD tasks at the levels of representation learning and anomaly discrimination (Mishra et al., 2024). Existing studies (Liu et al., 2018b, 2024d; Bao et al., 2022; Li et al., 2024a) have shown that high-performance VAD models need to target the modeling of appearance and motion information, i.e., the difference between regular events and anomalous examples in both spatial and temporal dimensions. In contrast to time series AD which mainly measures periodic temporal patterns of variables, and image AD which only focuses on spatial contextual deviations, VAD needs to extract both discriminative spatial and temporal features from a large amount of redundant information (e.g., repetitive temporal contexts and label-independent data shifts), as well as to learn the differences between normal and anomalous events in terms of the local appearances and global motions (Qiu et al., 2024; Liu et al., 2025c, 2022b).
@@
However, video anomalies are ambiguous and subjective (Liu et al., 2021b, 2023c). The same driving behavior can be classified differently depending on road conditions and contextual environments. For example, riding a horse in a grassland is usually normal, whereas a horse appearing on a highway would be considered an anomaly. On the one hand, compared to regular events, anomalies in the real world are difficult to be comprehensively predefined and have a much lower overall frequency of occurrence, making them difficult to collect. Labeling a sufficient number of all possible abnormal samples for model training is almost impossible. As a result, traditional supervised learning-based classification models are usually ineffective in dealing with AD tasks (Mašková et al., 2024). On the other hand, since video storage and transmission costs are significantly higher than other data modalities, engineers favor processing such data on the end or edge side to reduce communication overhead. As we all know, such devices, including surveillance cameras, smartphones, and local servers, are computationally resource-limited. Therefore, it is highly practical to develop deployable VAD systems for real-world applications, which requires concerted efforts by AI beyond communities.
@@
In this article, we extend the conventional scope of VAD from algorithm engineering on spatial-temporal anomaly detection to practical research towards real-world applications, termed Networking Systems for Video Anomaly Detection (NSVAD), to engage a broader readership from the IoT and computing communities. According to research objectives and involved domains, NSVAD is delineated into the hierarchical architecture shown in Fig. 1, encompassing: 1) Hardware Layer, consisting of various video sensors, communication units, and computing servers, etc, responsible for data acquisition, transmission, processing, and result reporting, as well as device networking; 2) System Layer, targeting resource optimization and algorithm deployment platforms for large-scale IoT applications, linking the Hardware Layer and the Algorithm Layer, supporting the configurable deployment of VAD tasks on various terminals; 3) Algorithm Layer, focusing on the development of detection algorithms and scene-specific models driven by artificial intelligence, especially deep learning; 4) Application Layer, encompassing IVSS in modern factories, agriculture, and smart cities, as well as various online video applications powered mobile internet. Most existing works belong to the Algorithm layer and solely concentrate on VAD model design, overlooking resource costs and challenges in real-world scenarios. For large-scale IoT and mobile video internet, NSVAD with stable performance and reasonable overheads that support online detection necessitates sensor networking research and support for resource optimization from computing communities. We review the recent advancements and typical methods in the algorithm layer and provide our latest explorations in the system layer to inspire readers to develop NSVAD toward real-world scenarios.
@@
Thanks to the development of edge AI (Walia et al., 2023) and artificial neural networks (Li et al., 2021b), deep learning-driven NSVAD algorithms have made significant progress in recent years and derived Unsupervised (UVAD), Weakly-supervised (WsVAD), and Fully-unsupervised (FuVAD) routes (Liu et al., 2024h). They have liberated human beings from massive videos analysis works and alleviated public information security concerns. Compared to early manual feature engineering, deep architectures such as convolutional neural networks and attention mechanisms can extract spatial-temporal representations from video sequences end-to-end without applying human a priori, empowering IVSS to process videos in different resolutions and scenarios. Therefore, researchers in this field are currently focusing on deep structure design and optimization learning strategies. They have creatively proposed multimodal VAD (Wu et al., 2020, 2022; Tian et al., 2021), Open-Set AVD (OSVAD) (Acsintoae et al., 2022; Zhu et al., 2022), Open-Vocabulary VAD (OVVAD) (Wu et al., 2024c), video anomaly segmentation (Tian et al., 2024), and anomaly retrieval (Wu et al., 2024b) tasks as well as integrated detection systems that can be deployed in practical scenes, such as modern manufacturing (Liu et al., 2024d), smart city (Liu et al., 2023f), and automated driving (Wang et al., 2021a). In addition to algorithm design, researchers from Networking Systems of AI (i.e., research on the deep convergence of communication and AI) (Song et al., 2022) and IoVT (i.e., subfields of IoT focusing on video sensor design, networking, and data processing) (Liu et al., 2021c) have begun to explore the design deployment-oriented VAD systems to collaboratively deal with multiple challenges (e.g., multi-view cross-scene heterogeneous videos and communication overheads) that come from the dynamic scenarios at the application layer and the limited resources at the hardware layer. These explorations and progress greatly expand the research boundaries and application scenarios of VAD, promoting it as an intelligent system science, i.e., NSVAD.
@@
Although there have been some reviews (Pang et al., 2021; Cook et al., 2019; Santhosh et al., 2020; Samaila et al., 2024) focusing on AD and combing its related work, due to the limited research horizon, the early works usually regard VAD as a fringe research task in the AD community. They focus on time series or images but lack an illuminating analysis of the AD task in video data. Recent survey papers (Nayak et al., 2021; Ramachandra et al., 2020; Chandrakala et al., 2022) continue to focus on unsupervised NSVAD routes in the same vein as the conventional AD task, i.e., using only normal samples to train generative models to learn the prototypical patterns of regular events and to discriminate uncharacterizable test samples as anomalies (Luo et al., 2021). Such reviews, while providing a comprehensive taxonomy of VAD research from the outlier detection perspective, have been informative in the last decade when unsupervised methods have dominated the NSVAD algorithm research. However, they ignore the emerging weakly-supervised (Feng et al., 2021a; Tian et al., 2021) and fully unsupervised (Pang et al., 2020; Zaheer et al., 2022) routes, which are of limited value in guiding further research.
@@
Considering the differences in knowledge bases and orientations of readers, this article provides an in-depth analysis of the basic concepts, related knowledge, and recent advances involved in NSVAD research and summarizes the available research resources. We systematically analyze the assumptions, frameworks, scenarios, advantages, and disadvantages of unsupervised, weakly supervised, and fully unsupervised VAD routes and explain in detail the relevant domain knowledge involved in each route. In addition, we introduce our NSVAD systems designed for dynamic environments in complex scenarios such as industrial IoT (Cao et al., 2025b) and smart cities (Liu et al., 2025b) to guide the NSVAD research in specific applications. Finally, we forecast the research challenges, trends, and possible opportunities to inspire future exploration.
@@ 1.1. Attention Analysis
We searched the number of publications and citations with the topic of ”video anomaly detection” in various mainstream academic databases (e.g., IEEE Xplore, ACM Digital Library, SpringerLink, ScienceDirect, and DBLP) to quantitatively present the research hotness, as shown in Fig. 2(a). Early VAD works were limited by hand-crafted features, which cannot handle complicated videos and rely on human a priori, thus having a slow start. Encouraged by the remarkable success of deep learning in video understanding tasks (e.g., action recognition, scene understanding, expression recognition (Wang et al., 2022a) and multimodal perception (Yang et al., 2022)), VAD research saw a boom after 2010, with an explosive growth in the number of publications and citations that continues to the present. On the one hand, the spread of surveillance cameras and streaming media platforms has provided sufficient data support for NSVAD research, making it possible to train large-scale deep models with high-performance GPUs. On the other hand, the increasing demand for offending video content detection in various scenarios drives many researchers and engineers from AI and IoT fields to devote themselves to NSVAD research.
@@ 1.2. Related Work
2021 (Nayak et al., 2021) Deep learning-based unsupervised VAD
@@
2021 (Pang et al., 2021) Deep learing-based unsupervised AD
@@
2024 (Liu et al., 2024h) Generalized VAD
2024 Ours NSVAD routes in video IoT
@@
To our knowledge, this is the first tutorial-type paper on NSVAD, providing a systematic overview of the basics, recent advances, and practical applications of various VAD routes. Previous papers (Liu et al., 2024h; Ramachandra et al., 2020; Chandrakala et al., 2022; Kiran et al., 2018) primarily focus on literature reviews, while conference tutorials lack systematic content, as shown in Table 1. Considering non-specialists’ limited background knowledge, this article emphasizes clear explanations of basic concepts and models. We introduce task definitions and learning frameworks for unsupervised, weakly supervised, and fully unsupervised VAD, as well as emerging tasks like open-set (Zhu et al., 2022), open vocabulary (Wu et al., 2024c), and glance VAD (Zhang et al., 2024b).
Initially, VAD was considered a fringe topic in the broader AD community and only briefly mentioned in AD overviews (Cook et al., 2019; Pang et al., 2021), lacking comprehensive surveys. Recent efforts have started organizing the state of VAD research (Ramachandra et al., 2020; Chandrakala et al., 2022; Raja et al., 2022; Nayak et al., 2024), but they focus mainly on unsupervised methods and overlook rising research routes like WsVAD (Sultani et al., 2018) and FuVAD (Wu et al., 2020). These routes offer reliable performance in real-world applications, and their importance is increasingly recognized. Fully unsupervised methods, in particular, allow efficient VAD model learning from large-scale video streams. Recent works, such as (Liu et al., 2024h), summarize these developments but do not cover emerging tasks like OSVAD (Acsintoae et al., 2022; Zhu et al., 2022) and OVVAD (Wu et al., 2024c), which are of crucial value for IoT applications.
Researchers in the AD community have paid attention to the continued progress and application prospects of VAD and introduced it to attendees through conference tutorials. For example, Pang et al. [^1] https://cvpr.thecvf.com/virtual/2023/tutorial/18560 organized a tutorial titled ”Recent Advances in Anomaly Detection” at CVPR2023, focusing on recent work on deep learning-driven unsupervised and weakly supervised VAD. We think this is a good start and highly appreciate the contribution of the organizers to the field. However, a more unified conceptual statement, more comprehensive documentation, and more systematic analysis of challenges and opportunities are necessary to inspire a wider community of readers.
Figure 3. Content navigation of this article.